# HA-VLN: A Benchmark for Human-Aware Navigation in Discrete–Continuous Environments with Dynamic Multi-Human Interactions, Real-World Validation, and an Open Leaderboard

### v1

@@ 7 Conclusion
We propose a Human-Aware VLN (HA-VLN) framework unifying discrete and continuous navigation under realistic multi-human conditions. By integrating dynamic human motion, refined annotations, and high-fidelity simulators, our HA-R2R dataset emphasizes human-centric instructions. Experiments show social awareness, multi-human interactions, and partial observability greatly increase complexity, reducing advanced agents’ performance. Nevertheless, our approach balances safety, efficiency, and personal space. Real-world tests confirm sim-to-real transfer, while our public leaderboard standardizes evaluations. By releasing all data, simulators, agents, and tools, we promote socially responsible, context-aware navigation in dynamic, human-populated environments.
@@
This supplementary material provides expanded details and results that complement the main paper. Section A offers a comprehensive literature survey focusing on three key research challenges. Section B describes our dataset construction, annotation protocols, real-time rendering methods, API design, and additional insights on annotation data. Section C presents an in-depth overview of the HA-R2R dataset and the proposed navigation agents. Finally, Section D includes detailed evaluation metrics, additional numerical results, visualized navigation outcomes, and real-world robot validation studies, each supplemented with thorough analysis. For further resources, the project page is accessible at https://ha-vln-project.vercel.app/, and the datasets are provided at DATASETS.
@@ Appendix A Related Work
This appendix surveys the evolution of Vision-and-Language Navigation (VLN) tasks, simulators, and agent designs, with particular attention to how Human-Aware VLN (HA-VLN) advances the state of the art. We focus on three key aspects deemed critical for bridging the Sim2Real gap: (1) Socially Compliant Navigation, (2) Human-Aligned Instructions and Visual Cues, and (3) Dynamic Environments with Human Activities and Interactions. Table A1 summarizes how prior work compares under these dimensions.
@@ A.1 Development of VLN Tasks
Early VLN tasks focused on basic indoor navigation—exemplified by Room-to-Room (R2R) [3, 15, 29]—and outdoor tasks like TOUCHDOWN [7] and MARCO [37]. Later efforts such as REVERIE [43] and VNLA [38] introduced object-centric or goal-driven navigation. While these approaches expanded the range of tasks, they typically overlooked real human behavior and social contexts. Dialogue-based tasks (e.g., DialFRED [13], CVDN [47]) incorporated interactive elements but did not account for dynamically moving bystanders or social-distance constraints. Initiatives like VLN-CE [28] moved closer to real-world conditions by enabling continuous (free-motion) navigation, yet remained largely devoid of explicit human factors [21, 29, 38, 47]. HA3D [31] addressed human motion and included human-oriented instructions but did not require agents to conform to social norms—e.g., maintaining safe distances or refraining from disturbing ongoing activities. Our proposed HA-VLN addresses these gaps by embedding all three essential elements—socially compliant navigation, human-referenced instructions, and dynamic human activities—into a single framework. Agents must plan routes among unpredictable bystanders, interpret language mentioning people and their behaviors, and uphold social standards. This integrated setup results in a benchmark that more closely aligns with real-world navigation demands.
@@ A.2 Simulators for VLN Tasks
A reliable simulator is essential for developing and evaluating VLN agents. Early simulators like Matterport3D [3] and House3D [52] offered photorealistic or synthetic indoor environments but lacked mobile humans. Others, such as AI2-THOR [26] and Gibson [53], introduced more interactive elements yet typically assumed static or purely synthetic contexts, thus limiting their applicability for studying social compliance. Google Street View, used in some outdoor navigation tasks, presents static imagery with occasional humans in the scene but lacks dynamic or interactive elements. HA3D [31] moved a step further by including human activities and instructions referencing people, though it did not mandate socially compliant navigation. HabiCrowd [48] integrated crowds into photorealistic domains, improving visual diversity but omitting human-aligned instructions. Similarly, Habitat 3.0 [45] provides high-performance simulation without extensive multi-human or social-compliance features.
By contrast, our HA-VLN Simulator unifies dynamic human activities, photorealistic rendering, and social-compliance requirements. Agents perceive and react to evolving bystander behaviors—such as avoiding collisions or maintaining personal space—using both discrete (panoramic) and continuous (free-motion) navigation. As shown in Table A2, HA-VLN surpasses prior simulators in terms of scene count, motion diversity, and instruction alignment, providing a comprehensive platform for training and testing socially aware VLN agents. Specifically, we introduce 675 scenes (across 90 scenarios), 122 motion types, and a cohesive framework that supports instruction-driven dynamic human interactions. By supporting both discrete and continuous action spaces, HA-VLN further broadens its potential for addressing diverse navigation goals and real-world deployment challenges.
Table A2: Comparison with Habitat 3.0 and HabiCrowd. Our environment extends beyond existing simulators by offering a larger variety of natural motion types, an increased number of scenario types, and explicit instruction alignment for human-aware navigation. “Num Scenes” and “Scenario Types” refer to the total annotated scans and the distinct activities or region annotations, respectively; “Natural Motions” indicates whether human motions are photorealistically rendered and not purely script-based; “Scene & Interaction” denotes whether dynamic multi-human interactions are included in the environment.
Simulator Habitat Series Num Scenes Scenario Types Natural Motions Motion Type Scene & Interaction Instr. Align
Habitat 3.0 [41] 3 59 59 Yes 2 Yes No
HabiCrowd [48] 2 480 No None No No
Ours 1 675 90 Yes 122 Yes Yes
@@ A.3 Agents for VLN Tasks
From early attention-based and reinforcement-learning approaches [36, 43, 50] to modern vision-language pre-training [35, 18, 32], VLN agents have grown increasingly adept at parsing instructions and navigating complex environments. However, most existing solutions—including EnvDrop [46], PREVALENT [18], and VLN-BERT [20]—rely on panoramic navigation, streamlining the action space but limiting realism of their movement. Recent efforts like NavGPT [57] and NaVid [55] explore continuous, egocentric navigation in partially dynamic worlds, yet they still lack explicit attention to human-aligned instructions or social compliance. In particular, these agents may not recognize the need to maintain safe distances, avoid disturbing activities, or adapt routes with active bystanders. Our HA-VLN agents address these gaps by navigating among multiple, moving humans and adhering to social norms. They interpret fine-grained, human-centric instructions and leverage visual cues that reflect real-world interactions, ensuring collision-free, respectful travel. This fusion of social compliance and dynamic human contexts sets HA-VLN apart, aligning agent behavior more closely with authentic, real-world challenges.
@@ Appendix B Simulator Details
We develop HAPS 2.0 to address the shortcomings of its predecessor, particularly in terms of mismatches between textual descriptions and motion data, as well as the limited diversity of region–motion associations.
@@
Altogether, HAPS 2.0 includes 26 distinct regions across 90 architectural scenes, covering 486 human activities in both indoor and outdoor contexts. Figure A2 illustrates these improvements. By offering more accurate, flexible, and diverse depictions of human actions, HAPS 2.0 provides a robust foundation for research in human motion analysis, social navigation, and beyond.
@@ B.2 Coarse Annotation Using PSO
We adopt a coarse-to-fine strategy for positioning human motions in 3D scans. Initially, we define each region by boundary coordinates $\mathbf{B}_{\text{lo}}=(x_{\text{lo}},y_{\text{lo}},z_{\text{lo}})$, $\mathbf{B}_{\text{hi}}=(x_{\text{hi}},y_{\text{hi}},z_{\text{hi}})$, and compile an object list $\mathbf{O}=\{j_{1},j_{2},\dots,j_{n}\}$ with positions $\mathbf{p}^{j_{i}}$. We then use Particle Swarm Optimization (PSO) [25] (Algorithm A1) to locate each motion $h_{i}$ at an optimal position $\mathbf{p}^{opt}$.
@@
Figure A2: Overview of HAPS 2.0 Scenes. These examples illustrate annotated human subjects across multiple scans in the HAPS 2.0 Dataset, highlighting a range of well-aligned motions, movements, and interactions (both with objects and with other humans).
@@ B.3 Fine Annotation Using a Multi-Camera Setup
To refine the coarse placements of human motions, we draw inspiration from 3D skeleton-capture methods [22, 40] and deploy nine RGB cameras, each positioned around the motion site. As shown in Figure A1, this arrangement provides a comprehensive multi-view perspective, revealing potential collisions or misalignments between the human figure and surrounding objects.
@@
$\tan\theta_{\text{ud}}^{i}=\begin{cases}0,&\text{if }i\text{ is odd},\\ \frac{\Delta_{z}}{\sqrt{2}\,\epsilon},&\text{if }i\text{ is even}.\end{cases}$
@@ B.4 Fine Annotation Protocol
    Initial View. Generate an overall preview of the human figure at $\mathbf{p}^{\text{h}}$ (Figure A1(a)).
@@
    Multi-Camera Observations. Collect images from the nine cameras (Figures A1(b)–(c)). Adjust camera angles or offsets as necessary, particularly in tight scenes like small bathrooms or closets.
@@ B.5 Multi-Human Interaction & Movement Enrichment
Human-in-the-Loop Method. We employ large language models (LLMs) such as ChatGPT-4 and LLaMA-3-8B-Instruct to propose plausible multi-human scenarios. Each prompt integrates details about existing human motions, object positions, and regional context, guiding the LLMs to generate rich, multi-character interactions. Our prompt design uses a system prompt and few-shot examples (Listings LABEL:list:dataset_prompt and LABEL:list:few-shot-example) to ensure clarity and detail. For instance, we collect each human’s position and identify objects within 6 m, describing relative distances and orientations. The LLMs then construct additional human activities suited to the scene, merging them into cohesive multi-person narratives.
Iterative Annotation Workflow. After the LLMs produce candidate interactions, we merge outputs from ChatGPT-4 and LLaMA-3-8B-Instruct, then manually refine and validate them over four rounds [11]. This process corrects inconsistencies and ensures contextual alignment. We subsequently place new human motions according to the generated descriptions, leveraging our multi-camera technique (Sec. B.3) for precise annotation of complex activities (e.g., stair-walking, see Fig. A3).
Examples of Enriched Interactions. Figure 2 demonstrates how additional humans can populate a living room: “two people sit on the couch, sharing popcorn on a small table,” while “a third friend stands in front of the TV, cheering.” Such enriched scenes capture realistic multi-human behaviors—from casual gatherings to active cheering—offering agents a broader range of social cues for navigation and interaction.
@@
2:  Initialize total signals sent and processed to 00;
@@
14:        Clear $\mathcal{Q}$ and reset total signals sent to 00;
@@ B.6 Real-Time Human Rendering
We integrate dynamic human models into the simulation through a multi-threaded pipeline inspired by Producer–Consumer principles and Java-style signaling (check details in Algorithm A2). This setup enables agents to observe and respond to human motions in real time, facilitating adaptable navigation policies.
@@
[…]
Synchronization & Consistency. We refresh human models immediately before the agent’s perception step, ensuring it observes the latest motion state. Upon starting a new episode, $\mathcal{Q}$ is cleared, and signal counters reset, so human motions revert to frame 0, maintaining consistency across episodes. This real-time process keeps human activities synchronized with the agent’s action cycle, creating dynamically evolving scenes where agents must adapt to changing bystander locations and behaviors.
Figure A4: The visualization of Human Counting.
Figure A5: Statistics on human motion annotations in HA-VLN simulator: (a) Distribution of humans by region; (b) Top 15 motions without multi-human interaction & movement enrichment; (c) Top 15 motions with enrichment; (d) Distribution of human trajectory lengths (in meters); (e) Impact of human presence on the environment, categorized as direct, indirect, and no impact. (Zoom in to view)
@@ B.7 API Design
Discrete Environment (DE). In our discrete setting, all agent and human positions are tracked via a real-time navigational graph displayed in a 2D top-down view. Each human’s activity is stored as a tuple $\langle p_{h},d_{agent},\theta_{relative},a_{status}\rangle$, where $p_{h}$ is the human’s 2D coordinate, $d_{agent}$ is the distance to the agent, $\theta_{relative}$ is the relative orientation, and $a_{status}$ indicates the activity state. This representation supports efficient, simultaneous tracking of multiple humans in a discrete viewpoint space.
@@
(1) Human Activity Monitoring. We track and analyze human activity in real time as in Sec. 3. When collisions occur, the agent reverts to its prior position, and we identify whether the obstacle is human or an inanimate object. For human collisions, we log the coordinates and motion state to inform potential reward-shaping strategies. Distance and orientation estimates derive from agent–human coordinate data. For instance, we employ the Grounding-DINO [34] detector on RGB inputs with the prompt “human” to count individuals. Figure A4 illustrates how human detection bounding boxes enable real-time counting.

### v3

@@ 6 Conclusion
We presented HA-VLN 2.0, a unified framework that standardizes discrete and continuous VLN under explicit human-centric constraints. By integrating dynamic human motion, refined annotations, and high-fidelity simulators, our HA-R2R dataset emphasizes human-centric instructions. Experiments show social awareness, multi-human interactions, and partial observability greatly increase complexity, reducing advanced agents’ performance. Nevertheless, our approach balances safety, efficiency, and personal space. Real-world tests confirm sim-to-real transfer, while our public leaderboard standardizes evaluations. By releasing all data, simulators, agents, and tools, we promote socially responsible, context-aware navigation in dynamic, human-populated environments.
@@
This supplementary material provides expanded details and results that complement the main paper. Section A offers a comprehensive literature survey focusing on three key research challenges. Section B describes our dataset construction, annotation protocols, real-time rendering methods, API design, and additional insights on annotation data. Section C presents an in-depth overview of the HA-R2R dataset and the proposed navigation agents. Finally, Section D includes detailed evaluation metrics, additional numerical results, visualized navigation outcomes, and real-world robot validation studies, each supplemented with thorough analysis. For further resources, access project page https://ha-vln-project.vercel.app/, and the datasets are provided at DATASETS.
@@ Appendix A Related Work
This appendix surveys the evolution of Vision-and-Language Navigation (VLN) tasks, simulators, and agent designs, with particular attention to how Human-Aware VLN (HA-VLN) 2.0 advances the state of the art. We focus on three key aspects deemed critical for bridging the Sim2Real gap: (1) Socially Compliant Navigation, (2) Human-Aligned Instructions and Visual Cues, and (3) Dynamic Environments with Human Activities and Interactions. Table A1 summarizes how prior work compares under these dimensions.
@@ A.1 Development of VLN Tasks
Early VLN tasks focused on basic indoor navigation—exemplified by Room-to-Room (R2R) Anderson et al. (2018); Fried et al. (2018); Gu et al. (2022); Ku et al. (2020)—and outdoor tasks like TOUCHDOWN Chen et al. (2019) and MARCO MacMahon et al. (2006). Later efforts such as REVERIE Qi et al. (2020) and VNLA Nguyen et al. (2019) introduced object-centric or goal-driven navigation. While these approaches expanded the range of tasks, they typically overlooked real human behavior and social contexts. Dialogue-based tasks (e.g., DialFRED Gao et al. (2022), CVDN Thomason et al. (2020)) incorporated interactive elements but did not account for dynamically moving bystanders or social-distance constraints. Initiatives like VLN-CE Krantz et al. (2020) moved closer to real-world conditions by enabling continuous navigation, yet remained devoid of explicit human factors Jain et al. (2019); Ku et al. (2020); Nguyen et al. (2019); Thomason et al. (2020). HA3D Li et al. (2024) addressed human motion and included human-oriented instructions but did not require agents to conform to social norms—e.g., maintaining safe distances or refraining from disturbing ongoing activities. Our proposed HA-VLN 2.0 addresses these gaps by embedding all three essential elements, socially compliant navigation, human-referenced instructions, and dynamic human activities, into a single framework. Agents must plan routes among unpredictable bystanders, interpret language mentioning people and their behaviors, and uphold social standards. This integrated setup results in a benchmark that closely aligns with real-world navigation demands.
@@ A.2 Simulators for VLN Tasks
A reliable simulator is essential for developing and evaluating VLN agents. Early simulators like Matterport3D Anderson et al. (2018) and House3D Wu et al. (2018) offered photorealistic or synthetic indoor environments but lacked mobile humans. Others, such as AI2-THOR Kolve et al. (2017) and Gibson Xia et al. (2018), introduced more interactive elements yet typically assumed static or purely synthetic contexts, thus limiting their applicability for studying social compliance. Google Street View, used in some outdoor navigation tasks, presents static imagery with occasional humans in the scene but lacks dynamic or interactive elements. HA3D Li et al. (2024) moved a step further by including human activities and instructions referencing people, though it did not mandate socially compliant navigation. HabiCrowd Vuong et al. (2024) integrated crowds into photorealistic domains, improving visual diversity but omitting human-aligned instructions. Similarly, Habitat 3.0 Savva et al. (2019) provides high-performance simulation without extensive multi-human or social-compliance features. By contrast, our HA-VLN Simulator unifies dynamic human activities, photorealistic rendering, and social-compliance requirements. Agents perceive and react to evolving bystander behaviors—such as avoiding collisions or maintaining personal space—using both discrete and continuous navigation. Specifically, we introduce 675 scenes (across 90 scenarios), 122 motion types, and a cohesive framework that supports instruction-driven dynamic human interactions. By supporting both discrete and continuous action spaces, HA-VLN further broadens its potential for addressing diverse navigation goals and real-world deployment challenges.
@@ A.3 Agents for VLN Tasks
From early attention-based and reinforcement-learning approaches Ma et al. (2019); Qi et al. (2020); Wang et al. (2019) to modern vision-language pre-training Lu et al. (2019); Hao et al. (2020); Li et al. (2020), VLN agents have grown increasingly adept at parsing instructions and navigating complex environments. However, most existing solutions, including EnvDrop Tan et al. (2019), PREVALENT Hao et al. (2020) and VLN-BERT Hong et al. (2021), rely on panoramic navigation, streamlining the action space but limiting realism of their movement. Recent efforts like NavGPT Zhou et al. (2024) and NaVid Zhang et al. (2024a) explore continuous, egocentric navigation in partially dynamic worlds, yet they still lack explicit attention to human-aligned instructions or social compliance. In particular, these agents may not recognize the need to maintain safe distances, avoid disturbing activities, or adapt routes with active bystanders. HA-VLN agents address these gaps by navigating among multiple, moving humans and adhering to social norms. They interpret fine-grained, human-centric instructions and leverage visual cues that reflect real-world interactions, ensuring collision-free, respectful travel. This fusion of social compliance and human dynamics sets HA-VLN apart, aligning agent behavior more closely with real-world challenges Dong et al. (2025).
Table A1: Comparison of VLN tasks, simulators, and agents based on (1) Socially Compliant Navigation, (2) Human-aligned Instructions and Visual Cues, and (3) Dynamic Environments with Human Activities.
Socially
Compliant
Navigation
Human-aligned
Instructions
and Visual Cues
Dynamic
Environments
Prior Work
Tasks × × × MARCO MacMahon et al. (2006), DRIF Blukis et al. (2018), VLN-R2R Anderson et al. (2018), TOUCHDOWN Chen et al. (2019), REVERIE Qi et al. (2020), DialFRED Gao et al. (2022)
VNLA Nguyen et al. (2019), CVDN Thomason et al. (2020), R4R Jain et al. (2019), RxR Ku et al. (2020), EQA Das et al. (2018), IQA Gordon et al. (2018)
× × VLN-CE Krantz et al. (2020)
× HA3D Li et al. (2024)
HA-VLN (Ours)
Simulators × × × Matterport3D Anderson et al. (2018), House3D Wu et al. (2018), AI2-THOR Kolve et al. (2017), Gibson GANI Xia et al. (2018)
× × Habitat Savva et al. (2019), Google Street, ViZDoom Kempka et al. (2016)
× HA3D Li et al. (2024)
HA-VLN (Ours), Habitat3.0 Puig et al. (2023)
Agents × × × EnvDrop Tan et al. (2019), AuxRN Zhu et al. (2020), PREVALENT Hao et al. (2020), RelGraph Hong et al. (2020), HAMT Chen et al. (2021), NavCoT Lin et al. (2025)
Rec-VLNBERT Hong et al. (2021), EnvEdit Li et al. (2022), Airbert Guhur et al. (2021), Lily Lin et al. (2023), ScaleVLN Wang et al. (2023)
× NavGPT Zhou et al. (2024), NaVid Zhang et al. (2024a), Student Force Anderson et al. (2018)
HA-VLN Agent (Ours)
@@ Appendix B Simulator Details
We develop HAPS 2.0 to address the shortcomings of its predecessor Li et al. (2024), particularly in terms of mismatches between textual descriptions and motion data, as well as the limited diversity of region–motion associations.
@@
Altogether, HAPS 2.0 includes 26 distinct regions across 90 architectural scenes, covering 486 human activities in both indoor and outdoor contexts. Fig. A2 illustrates these improvements. By offering more accurate, flexible, and diverse depictions of human actions, HAPS 2.0 provides a robust foundation for research in human motion analysis, social navigation, and beyond.
@@ B.2 Coarse Annotation Using PSO
We adopt a coarse-to-fine strategy for positioning human motions in 3D scans. Initially, we define each region by boundary coordinates $\mathbf{B}_{\text{lo}}=(x_{\text{lo}},y_{\text{lo}},z_{\text{lo}})$, $\mathbf{B}_{\text{hi}}=(x_{\text{hi}},y_{\text{hi}},z_{\text{hi}})$, and compile an object list $\mathbf{O}=\{j_{1},j_{2},\dots,j_{n}\}$ with positions $\mathbf{p}^{j_{i}}$. We then use Particle Swarm Optimization (PSO) Kennedy & Eberhart (1995) (more details are provided in Algorithm A1) to locate each motion $h_{i}$ at an optimal position $\mathbf{p}^{opt}$.
@@
Figure A2: Overview of HA-VLN Scenes. These examples illustrate annotated human subjects across multiple scans in the HA-VLN simulator, highlighting a range of well-aligned motions, movements, and interactions (both with objects and with other humans).
@@ B.3 Fine Annotation Using a Multi-Camera Setup
To refine the coarse placements of human motions, we draw inspiration from 3D skeleton-capture methods Ji et al. (2018); Petrovich et al. (2021) and deploy nine RGB cameras, each positioned around the motion site. As shown in Fig. A1, this arrangement provides a comprehensive multi-view perspective, revealing potential collisions or misalignments between the human figure and surrounding objects.
@@
$\tan\theta_{\text{ud}}^{i}=\begin{cases}0,&\text{if }i\text{ is odd},\\ \frac{\Delta_{z}}{\sqrt{2}\,\epsilon},&\text{if }i\text{ is even}.\end{cases}$ (A1)
@@ B.4 Fine Annotation Protocol
    Initial View. Generate an overall preview of the human figure at $\mathbf{p}^{\text{h}}$ (Fig. A1(a)).
@@
    Multi-Camera Observations. Collect images from the nine cameras (Figs. A1(b)–(c)). Adjust camera angles or offsets as necessary, particularly in tight scenes like small bathrooms or closets.
@@ B.5 Multi-Human Interaction & Movement Enrichment
Human-in-the-Loop Method. We employ large language models (LLMs) such as ChatGPT-4 and LLaMA-3-8B-Instruct to propose plausible multi-human scenarios. Each prompt integrates details about existing human motions, object positions, and regional context, guiding the LLMs to generate rich, multi-character interactions. Our prompt design uses a system prompt and few-shot examples (Listings 1 and 2) to ensure clarity and detail. For instance, we collect each human’s position and identify objects within 6 m, describing relative distances and orientations. The LLMs then construct additional human activities suited to the scene, merging them into cohesive multi-person narratives.
Iterative Annotation Workflow. After the LLMs produce candidate interactions, we merge outputs from ChatGPT-4 and LLaMA-3-8B-Instruct, then manually refine and validate them over four rounds Ding et al. (2024); Cheng et al. (2024). This process corrects inconsistencies and ensures contextual alignment. We subsequently place new human motions according to the generated descriptions, leveraging our multi-camera technique (Sec. B.3) for precise annotation of complex activities (e.g., stair-walking, see Fig. A3).
Examples of Enriched Interactions. Fig. 2 demonstrates how additional humans can populate a living room: “two people sit on the couch, sharing popcorn on a small table,” while “a third friend stands in front of the TV, cheering.” Such enriched scenes capture realistic multi-human behaviors—from casual gatherings to active cheering—offering agents a broader range of social cues for navigation and interaction.
@@
2: Initialize total signals sent and processed to 0;
@@
14:   Clear $\mathcal{Q}$ and reset total signals sent to 0;
@@ B.6 Real-Time Human Rendering
We integrate dynamic human models into simulation through a multi-threaded pipeline inspired by Producer–Consumer principles and Java-style signaling (Algorithm A2). This setup enables agents to observe and respond to human motions in real time, facilitating adaptable navigation policies.
@@
Main Simulation Thread (Thread 2). When the agent is about to act, Thread 2 checks $\mathcal{Q}$ for pending refresh signals. It calculates the current frame index $t$ as $\text{(signals\_processed}\bmod N)$, where $N$ is the total length of the human motion sequence. Template manager $\mathcal{T}$ then removes outdated models and loads frame $t$ into the environment, adjusting each figure’s position and orientation.
Synchronization & Consistency. We refresh human models immediately before the agent’s perception step, ensuring it observes the latest motion state. Upon starting a new episode, $\mathcal{Q}$ is cleared, and signal counters reset, so human motions revert to frame 0, maintaining consistency across episodes. This real-time process keeps human activities synchronized with agent’s action cycle, creating dynamic scenes where agents must adapt to changing bystander locations and behaviors.
Figure A4: Statistics on human annotations in HA-VLN simulator: (a) Distribution of humans by 26 region types; (b) Top 15 motions without multi-human interaction & movement enrichment; (c) Top 15 motions with enrichment; (d) Distribution of human trajectory lengths (in meters); (e) Impact of human presence on environment, categorized as direct, indirect, and no impact. (Zoom in to view)
@@ B.7 API Design
Discrete Environment (DE). In our discrete setting, all agent and human positions are tracked via a real-time navigational graph displayed in a 2D top-down view. Each human’s activity is stored as a tuple $\langle p_{h},d_{agent},\theta_{relative},a_{status}\rangle$, where $p_{h}$ is the human’s 2D coordinate, $d_{agent}$ is the distance to the agent, $\theta_{relative}$ is the relative orientation, and $a_{status}$ indicates activity state. This representation supports efficient, simultaneous tracking of multiple humans in a discrete viewpoint space.
@@
Figure A5: The visualization of Human Counting.
(1) Human Activity Monitoring. We track and analyze human activity in real time as in Sec. 3. When collisions occur, the agent reverts to its prior position, and we identify whether the obstacle is human or an inanimate object. For human collisions, we log the coordinates and motion state to inform potential reward-shaping strategies. Distance and orientation estimates derive from agent–human coordinate data. For instance, we employ the Grounding-DINO Liu et al. (2024) detector on RGB inputs with the prompt “human” to count individuals. Fig. A5 illustrates how human detection bounding boxes enable real-time counting.